Our operations team will perform an emergency maintenance on storage units today, Monday 9 June 2014.

This maintenance may impact the following services :
  • www.gandi.net: access to the control panel, orders
  • mail.gandi.net: IMAP, POP, SMTP, webmail
  • Web redirections
  • GandiSite (Basekit) / Sitemaker
  • DNS zone provisioning
  • Gandi Groups / Gandi wiki
  • Emergency console / DNS cache for hosting datacenter in Paris
  • SimpleHosting (PaaS) virtualhost provisioning

The maintenance window will be four hours, and will occur between 2 and 4pm PDT (11pm to 3am CEST/Paris).

We apologize for the inconvenience this maintenance may cause.

 

Updates:

2:00pm PDT [11:00 PM CEST, 21:00 UTC]: start of the maintenance

2:47pm PDT: end of the maintenance


A scheduled maintenance will take place on Wednesday 26/02 22:00 - 00:00 (UTC)

During the maintenance, zone file updates on domains hosted on {a,b,c}.dns.gandi.net could be delayed.

Please excuse the inconveniance.

Regards;

 

Gandi Team


A scheduled maintenance will take place on Sunday 23/02 22:00 - 23:00 (UTC)

During the maintenance, zone file updates on domains hosted on {a,b,c}.dns.:gandi.net coiuld be delayed.

Please excuse us for any inconveniance.

Regards;

 

Gandi Team


Our database system will have a maintenance tonight.
Therefore all update of the DNS zones updates will be delayed betweeen 00:00 CEST and 02:00 CEST.
We apologize for the inconvenience.

We will be performing major network upgrades to our backbone network in Paris during the night of 23-24 March 2013.  

 

The purpose of this work is to replace our current (aging) backbone network routers with newer generation equipment allowing for increased capacity, performance and resilience.  During this time, all Gandi services in Paris will experience brief periods of network unavailability and/or instability as the migration proceeds.

 

This activity will not impact the individual Gandi services themselves, but rather only the network connectivity between the internet and those services.  As a result, there will be no need for hosting customers to restart their IaaS or PaaS servers/instances.  Services in our Baltimore datacentre will not be directly affected by this maintenance, but some periods of routing instability and reconvergence may be experienced, though these will be far less prevalent than for the Paris services.

 

This migration will commence at 01:30 CET (00:30 GMT) on Sunday 24 March 2013, and we estimate between 10 and 15 minutes of major network outage during the initial migration, followed by a number of brief periods of network instability during the remainder of the migration.  The entire migration activity should last approximately 90 minutes, but no more than two hours.

Operations will be ended at 05:00 AM CET (04:00 AM GMT).

 

We apologise in advance for any inconvenience this may cause.

EDIT 03:08 AM CET : maintenance ended


During the night of Tuesday the 19th to Wednesday the 20th of June 2012, we are going to interrupt the generation of DNS zones in order to perform a database modification.

This modification will notably allow inserting long records (SPF or TXT for example).

Modifications made during this time period will therefore be put on hold and will be applied only once our operation is over.

Sincerely,
The technical team

[Resolved] This maintenance was resolved without incident ]



We have lost contact with our infrastructure at Telehouse1 (Jeuneurs) in Paris as of 03:11 CET (02:11 GMT).  Our technicians are en-route to investigate the problem.  

 

Affected Gandi services are:

1/3 of the DNS systems, whois for domain registry, and mail redirection service.

 

We will update here as we obtain more details.

 

UPDATE: 04:01 CET (03:01 GMT) - our core router at Telehouse1 had suffered a system crash resulting in the outage of this site.  Our engineer successfully restarted the router and we have obtained crash dump information from the router for analysis.  All services are once again operational.


Due to a technical issue, DNS zone updates are not being processed since midday.  Our technical teams are investigating the cause of the problem and are working to resolve the issue as soon as possible.

 

Update: 14:00 CET :  The problem has been resolved.


Page 1 2
Change the news ticker size